Shaver Co. manufactures a variety of electric razors for men and women. The company's plant is partially automated. Listed below is cost driver information used in the product-costing system:

Overhead Cost Pool Budgeted Overhead Cost Budgeted Cost Driver Level Cost Driver
Machinery depreciation/maintenance $ 199,822 28,546 Machine hours
Factory depreciation/utilities/insurance 142,730 28,546 Machine hours
Product design 576,000 42,980 Hours in design
Material handling 1,624,000 203,000 Pounds of raw materials

In addition, Shaver expects to spend $570,920 for 8,156 direct labor-hours. Two current product orders had the following requirements:

Men’s Razors Women’s Razors
Units produced and sold 21,000 26,000
Direct labor hours 20 30
Pounds of raw materials 793 683
Hours in design 10 13
Machine hours 65 50

Using ABC, how much batch-level overhead is assigned to the current order for Women's Razors based on pounds of raw materials?

Multiple Choice

  • $5,464.

  • $5,687.

  • $5,157.

  • $16,307.

  • $6,309.
Answer #1

Predetermined overhead rate = 1624000/203000 = $8 per pound of raw material

Overhead assigned to the current order for women razors = 8*683

= $5464